Prove that. In ∆abc Tan(a-b-c)=tan2a ....what is ans??

Grade:11

1 Answers

Ravi
6 years ago
In a triangle a+b+c=180,so c+b=180-a ,put it in LHS Tan(a-(180-a))=Tan (2a-180) which is equal to Tan(2a)-Tan(180)/1+Tan(2a)*Tan(180)=Tan(2a)  because Tan(180)=0,
 Lhs=Rhs
Hence proved